an airplane flies 500km against the wind in two hours. it flies the same distance with the wind in one hour. what is the airplane's speed in still air?
Answer by lwsshak3(11628)

an airplane flies 500km against the wind in two hours. it flies the same distance with the wind in one hour. what is the airplane's speed in still air?
***
let x=speed of plane in still air
let c=speed of wind
x+c=speed of plane with the wind
x-c=speed of plane against the wind
distance =speed*travel time
..
in 1 hr, plane travels 250 km against the wind
in 1 hr, plane travels 500 km with the wind
..
250=(x-c)(1)
500=(x+c)(1)
add
750=2x
x=375
c=x-250=125
speed of plane in still air=375 km/hr